# Initial value space of the fourth order Painlevé system with $A_{5}^{(1)}\times A_{1}^{(1)}$ symmetry

### v2

@@ 4.1 Blowing up
The relation between the coordinates $\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)}\right)$ and $\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)$ is given by
$q_{1}=1+\left(\alpha_{0}-x_{2}^{[01]}z_{2}^{[01]}\right)z_{2}^{[01]},\quad q_{2}=y_{2}^{[01]},\quad p_{1}^{(01)}=z_{2}^{[01]},\quad p_{2}^{(01)}=z_{2}^{[01]}w_{2}^{[01]}.$ (4.6)
Then we see that the accessible singularity $C_{2}^{(01)}$ corresponds to $\left\{z_{2}^{[02]}=0\right\}$. Hence all of the solutions to system (4.1) passing through $C_{2}^{(01)}$ can be parametrized in terms of an initial condition
$\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)=(a_{0},b_{0},0,d_{0}),$ (4.7)
where $a_{0},b_{0},d_{0}\in\mathbb{C}$, for system (4.5). Combining relations (2.4) and (4.6), we obtain
$x_{2}^{[01]}=-\{(q_{1}-1)p_{1}-\alpha_{0}\}p_{1},\quad y_{2}^{[01]}=q_{2},\quad z_{2}^{[01]}=\dfrac{1}{p_{1}},\quad w_{2}^{[01]}=p_{2}$ (4.8)
@@
and it is a symplectic transformation. Moreover, system (4.5) is described as a polynomial Hamiltonian system. We can show it by a direct calculation.
@@ 4.2 Formal power series
$\left(q_{1}(t_{0}),q_{2}(t_{0}),p_{1}^{(01)}(t_{0}),p_{2}^{(01)}(t_{0})\right)=(1,b_{0},0,0),$
where $b_{0}\in\mathbb{C}$. This solution is expressed in terms of formal power series as
$\begin{split}&q_{1}=1+(t-t_{0})^{\lambda_{a}}\sum_{k=1}^{\infty}a_{k}(t-t_{0})^{k},\quad q_{2}=b_{0}+(t-t_{0})^{\lambda_{b}}\sum_{k=1}^{\infty}b_{k}(t-t_{0})^{k},\\ &p_{1}^{(01)}=(t-t_{0})^{\lambda_{c}}\sum_{k=1}^{\infty}c_{k}(t-t_{0})^{k},\quad p_{2}^{(01)}=(t-t_{0})^{\lambda_{d}}\sum_{k=1}^{\infty}d_{k}(t-t_{0})^{k},\end{split}$ (4.9)
@@
$t_{0}(t_{0}-1)(\lambda_{c}+1)c_{1}(t-t_{0})^{\lambda_{c}}+O\left((t-t_{0})^{\lambda_{c}+1}\right)=-t_{0}+1+o(1).$
Here we denote the function $f(t)$ satisfying $\lim_{t\to t_{0}}\left|\frac{f(t)}{(t-t_{0})^{k}}\right|<\infty$ and $\lim_{t\to t_{0}}\left|\frac{f(t)}{(t-t_{0})^{k}}\right|=0$ by $O\left((t-t_{0})^{k}\right)$ and $o\left((t-t_{0})^{k}\right)$ respectively. It follows that
$\lambda_{c}=0,\quad c_{1}=-\frac{1}{t_{0}}.$ (4.10)
@@
$\lambda_{a}=0,\quad a_{1}=-\frac{\alpha_{0}}{t_{0}}$ (4.11)
from the first equation of system (4.1),
$\lambda_{b}=0$ (4.12)
from the second equation and
$\lambda_{d}=0$ (4.13)
from the fourth equation. Under conditions (4.10), (4.11), (4.12) and (4.13), we also obtain
$b_{1}=-\frac{2b_{0}(b_{0}-1)(b_{0}-t_{0})}{t_{0}-1}d_{1}+B_{1},$
where $B_{1}\in\mathbb{C}(t_{0})[\alpha_{0},\ldots,\alpha_{5},\eta,b_{0}]$, from the second equation. Then we see that $d_{1}$ turns to be an arbitrary complex number and $b_{1}$ is determined uniquely by $d_{1}$. Moreover, in a similar manner, $c_{2}$ and $d_{2}$ are determined uniquely as elements of $\mathbb{C}(t_{0})[\alpha_{0},\ldots,\alpha_{5},\eta,b_{0},d_{1}]$ and we obtain
$b_{2}=-\frac{(b_{0}+1)(b_{0}-t_{0})}{2(t_{0}-1)}a_{2}+B_{2},$
where $B_{2}\in\mathbb{C}(t_{0})[\alpha_{0},\ldots,\alpha_{5},\eta,b_{0},d_{1}]$. Then we see that $a_{2}$ turns to be an arbitrary complex number and $b_{2}$ is determined uniquely by $a_{2}$.
@@
$\frac{q_{1}-1}{p_{1}^{(01)}}$ $=\frac{-\alpha_{0}t_{0}^{-1}(t-t_{0})+a_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right)}{-t_{0}^{-1}(t-t_{0})+c_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right)}$
$=\frac{-\alpha_{0}t_{0}^{-1}+a_{2}(t-t_{0})+O\left((t-t_{0})^{2}\right)}{-t_{0}^{-1}+c_{2}(t-t_{0})+O\left((t-t_{0})^{2}\right)}$
$=\alpha_{0}+\frac{-t_{0}^{-1}a_{2}+\alpha_{0}t_{0}^{-1}c_{2}}{t_{0}^{-2}}(t-t_{0})+O\left((t-t_{0})^{2}\right)$
$=\alpha_{0}+t_{0}(-a_{2}+\alpha_{0}c_{2})(t-t_{0})+O\left((t-t_{0})^{2}\right),$
@@
$\frac{\frac{q_{1}-1}{p_{1}^{(01)}}-\alpha_{0}}{p_{1}^{(01)}}$ $=\frac{t_{0}(-a_{2}+\alpha_{0}c_{2})(t-t_{0})+O\left((t-t_{0})^{2}\right)}{-t_{0}^{-1}(t-t_{0})+O\left((t-t_{0})^{2}\right)}$
$=t_{0}^{2}(a_{2}-\alpha_{0}c_{2})+O(t-t_{0}).$
We also have
$\frac{p_{2}^{(01)}}{p_{1}^{(01)}}=-t_{0}d_{1}+O(t-t_{0})$
in a similar manner. Set
$x_{2}^{[01]}=-\frac{\frac{q_{1}-1}{p_{1}^{(01)}}-\alpha_{0}}{p_{1}^{(01)}},\quad y_{2}^{[01]}=q_{2},\quad z_{2}^{[01]}=p_{1}^{(01)},\quad w_{2}^{[01]}=\frac{p_{2}^{(01)}}{p_{1}^{(01)}}.$ (4.14)
@@
$\left(x_{2}^{[01]}(t_{0}),y_{2}^{[01]}(t_{0}),z_{2}^{[01]}(t_{0}),w_{2}^{[01]}(t_{0})\right)=\left(t_{0}^{2}(a_{2}-\alpha_{0}c_{2}),b_{0},0,-t_{0}d_{1}\right),$
which is essentially equivalent to (4.7). Note that relation (4.14) is equivalent to (4.6).
5 Resolving the accessible singularities $P_{j}^{(kl)}$
It is not suitable to use the blowing up method to resolve the accessible singularities (3.4), because the explicit formula of the system after blowing up becomes more complicated. Therefore we use the method in terms of formal power series and the symmetry of system (1.1) in this section. We give the result of this section in advance.

### v3

@@ 5.1 Blowing up
where $X_{3},Y_{3},Z_{3},W_{3}\in\mathbb{C}\left[x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]},t\right]$, which does not have any singularity on $D_{1}^{2}$. Then the relation between the coordinates $\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)}\right)$ and $\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)$ is given by
$q_{1}=-\left(x_{3}^{[01]}z_{3}^{[01]}+y_{3}^{[01]}w_{3}^{[01]}-\alpha_{5}+\eta\right)z_{3}^{[01]},\quad q_{2}=y_{3}^{[01]}z_{3}^{[01]},\quad p_{1}^{(01)}=z_{3}^{[01]},\quad p_{2}^{(01)}=w_{3}^{[01]}.$ (5.11)
Note that the accessible singularity $C_{3}^{(01)}$ corresponds to $\left\{z_{3}^{[01]}=0\right\}$. Combining relations (3.4) and (5.11), we obtain
$x_{3}^{[01]}=-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1},\quad y_{3}^{[01]}=q_{2}p_{1},\quad z_{3}^{[01]}=\dfrac{1}{p_{1}},\quad w_{3}^{[01]}=\frac{p_{2}}{p_{1}}$ (5.12)
@@
and it is a symplectic transformation. Moreover, system (5.10) is described as a polynomial Hamiltonian system. We can show it by a direct calculation. Note that all of the leaves corresponding to the solutions to system (5.10) passing through the set $\left\{z_{3}^{[01]}=0\right\}$ are transversal leaves. Besides, system (5.3) on the coordinate $\left(x^{\prime}_{31},y^{\prime}_{31},z^{\prime}_{31},w^{\prime}_{31}\right)$ is described as
$t(t-1)z^{\prime}_{31}dx^{\prime}_{31}-X^{\prime}_{31}dt=0,$
$t(t-1)z^{\prime}_{31}dy^{\prime}_{31}-Y^{\prime}_{31}dt=0,$
$t(t-1)dz^{\prime}_{31}-Z^{\prime}_{31}dt=0,$
$t(t-1)dw^{\prime}_{31}-W^{\prime}_{31}dt=0,$
where $X^{\prime}_{31},Y^{\prime}_{31},Z^{\prime}_{31},W^{\prime}_{31}\in\mathbb{C}\left[x^{\prime}_{31},y^{\prime}_{31},z^{\prime}_{31},w^{\prime}_{31},t\right]$. The singularity of this system on $D_{1}^{2}$ is given by
$\left\{x^{\prime}_{31}=z^{\prime}_{31}=0\right\}.$
We extend the system $\mathcal{H}$ to that on $\left(\Sigma_{\eta}\cup D_{1}^{1}\cup D_{1}^{2}\right)\times B$ and denote it by $\mathcal{H}_{1}^{2}$ temporarily. Then all of the leaves corresponding to the solutions to $\mathcal{H}_{1}^{2}$ passing through the singularity $\left\{x^{\prime}_{31}=z^{\prime}_{31}=0\right\}$ are vertical leaves. This fact can be shown in a similar manner as in $D^{0}\setminus(C^{0}\cup P^{0})$. We can consider the other accessible singularities similarly and obtain
$x_{1}^{[01]}=-\{(q_{1}-q_{2})p_{1}-\alpha_{2}\}p_{1},\quad y_{1}^{[01]}=q_{2},\quad z_{1}^{[01]}=\dfrac{1}{p_{1}},\quad w_{1}^{[01]}=p_{1}+p_{2},$
$x_{3}^{[02]}=q_{1}p_{2},\quad y_{3}^{[02]}=-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{2},\quad z_{3}^{[02]}=\dfrac{p_{1}}{p_{2}},\quad w_{3}^{[02]}=\dfrac{1}{p_{2}},$
$x_{2}^{[12]}=q_{1}^{(1)},\quad y_{2}^{[12]}=-\left\{\left(q_{2}^{(1)}-1\right)p_{2}^{(1)}-\alpha_{2}\right\}p_{2}^{(1)},\quad z_{2}^{[12]}=p_{1}^{(1)},\quad w_{2}^{[12]}=\dfrac{1}{p_{2}^{(1)}}.$
In the last, we consider the accessible singularity $P_{1}^{1}$. There exist two transversal leaves corresponding to solutions to system (5.5) passing through $P_{1}^{1}$ at any $t=t_{0}$. They are expressed in terms of formal power series as
$u_{12}^{\prime}=\frac{\alpha_{2}-\alpha_{5}+\eta}{t_{0}-1}(t-t_{0})+a_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right),$
$u_{22}^{\prime}=1-\frac{\alpha_{2}}{2t_{0}}+b_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right),$
$v_{12}^{\prime}=\frac{1}{2t_{0}}(t-t_{0})+c_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right),$
$v_{22}^{\prime}=-1+d_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right)$
and
$u_{12}^{\prime}=-\frac{\alpha_{2}-\alpha_{5}+\eta}{t_{0}-1}(t-t_{0})+a_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right),$
$u_{22}^{\prime}=1-\frac{\alpha_{2}}{t_{0}}+b_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right),$
$v_{12}^{\prime}=\frac{1}{t_{0}}(t-t_{0})+c_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right),$
$v_{22}^{\prime}=-1-\frac{\alpha_{2}-\alpha_{5}+\eta}{t_{0}}(t-t_{0})+d_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right).$
Here we denote the function $f(t)$ satisfying $\lim_{t\to t_{0}}\left|\frac{f(t)}{(t-t_{0})^{k}}\right|<\infty$ by $O\left((t-t_{0})^{k}\right)$. Two coefficients among $a_{2},b_{2},c_{2},d_{2}$ are not determined and turn to be arbitrary complex numbers. Since the former solutions passes through the accessible singularity $C_{1}^{(01)}$, we have to give coordinates which separate the latter solutions. By using the method which will be given in Section 6, we obtain two coordinates
$\left(\widehat{x}_{3}^{[01]},\widehat{y}_{3}^{[01]},\widehat{z}_{3}^{[01]},\widehat{w}_{3}^{[01]}\right)=r_{2}\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right),$
$\left(\widehat{x}_{3}^{[02]},\widehat{y}_{3}^{[02]},\widehat{z}_{3}^{[02]},\widehat{w}_{3}^{[02]}\right)=r_{2}\left(x_{3}^{[02]},y_{3}^{[02]},z_{3}^{[02]},w_{3}^{[02]}\right).$
We do not state its detail here.
@@ 5.2 Formal power series
We consider $C_{2}^{(12)}$ of $C_{1}^{0}$ in this subsection. We can prove for $C_{1}^{(01)}$, $C_{3}^{(01)}$ and $C_{3}^{(02)}$ in a similar manner. The system $\mathcal{H}$ on $W_{12}\times B$ is described as
$\begin{split}&t(t-1)p_{2}^{(12)}dq_{1}^{(1)}-F_{1}^{(12)}dt=0,\\ &t(t-1)p_{2}^{(12)}dq_{2}^{(1)}-F_{2}^{(12)}dt=0,\\ &t(t-1)p_{2}^{(12)}dp_{1}^{(12)}-G_{1}^{(12)}dt=0,\\ &t(t-1)dp_{2}^{(12)}-G_{2}^{(12)}dt=0,\end{split}$ (5.13)
where
$F_{1}^{(12)}$ $=\left[(\alpha_{5}+\eta)t\left(q_{1}^{(1)}\right)^{2}+\alpha_{3}q_{1}^{(1)}q_{2}^{(1)}-\left\{(\alpha_{0}+\alpha_{5}+\eta)t-\alpha_{0}-\alpha_{1}+\eta\right\}q_{1}^{(1)}-\alpha_{3}q_{2}^{(1)}-\alpha_{1}+\eta\right]p_{2}^{(12)}$
$\quad+2q_{1}^{(1)}\left(q_{1}^{(1)}-1\right)\left(tq_{1}^{(1)}-1\right)p_{1}^{(12)}+\left(q_{1}^{(1)}-1\right)\left(tq_{1}^{(1)}+q_{2}^{(1)}\right)\left(q_{2}^{(1)}-1\right),$
$F_{2}^{(12)}$ $=\left[\alpha_{5}tq_{1}^{(1)}q_{2}^{(1)}+(\alpha_{3}+\eta)\left(q_{2}^{(1)}\right)^{2}-\alpha_{5}tq_{1}^{(1)}+\{(\alpha_{1}+\alpha_{2}-\eta)t-\alpha_{2}-\alpha_{3}-\eta\}q_{2}^{(1)}-(\alpha_{1}-\eta)t\right]p_{2}^{(12)}$
$\quad+\left(q_{1}^{(1)}-1\right)\left(tq_{1}^{(1)}+q_{2}^{(1)}\right)\left(q_{2}^{(1)}-1\right)p_{1}^{(12)}+2q_{2}^{(1)}\left(q_{2}^{(1)}-1\right)\left(q_{2}^{(1)}-t\right),$
$G_{1}^{(12)}$ $=\eta\left(\alpha_{3}p_{1}^{(12)}-\alpha_{5}t\right)\left(p_{2}^{(12)}\right)^{2}+\Big[\alpha_{3}\left(q_{1}^{(1)}-1\right)\left(p_{1}^{(12)}\right)^{2}$
$\quad+\left\{(\alpha_{3}+2\eta)q_{2}^{(1)}-(\alpha_{5}+2\eta)tq_{1}^{(1)}+(-\alpha_{3}-\alpha_{4}+1)t+\alpha_{4}+\alpha_{5}-1\right\}p_{1}^{(12)}-\alpha_{5}tq_{2}^{(1)}+\alpha_{5}t\Big]p_{2}^{(12)}$
$\quad-\left\{2t\left(q_{1}^{(1)}\right)^{2}-2q_{1}^{(1)}q_{2}^{(1)}-(t+1)q_{1}^{(1)}+2q_{2}^{(1)}\right\}\left(p_{1}^{(12)}\right)^{2}$
$\quad-\left\{2tq_{1}^{(1)}q_{2}^{(1)}-2\left(q_{2}^{(1)}\right)^{2}-2tq_{1}^{(1)}+(t+1)q_{2}^{(1)}\right\}p_{1}^{(12)},$
$G_{2}^{(12)}$ $=\alpha_{3}\eta\left(p_{2}^{(12)}\right)^{2}$
$\quad+\left[\alpha_{3}\left(q_{1}^{(1)}-1\right)p_{1}^{(12)}+\alpha_{5}tq_{1}^{(1)}+2(\alpha_{3}+\eta)q_{2}^{(1)}+\left(\alpha_{1}+\alpha_{2}-\eta\right)t-\alpha_{2}-\alpha_{3}-\eta\right]p_{2}^{(12)}$
$\quad+\left(q_{1}^{(1)}-1\right)\left(tq_{1}^{(1)}+2q_{2}^{(1)}-1\right)p_{1}^{(12)}+3\left(q_{2}^{(1)}\right)^{2}-2(t+1)q_{2}^{(1)}+t.$
@@
$\left(q_{1}^{(1)}(t_{0}),q_{2}^{(1)}(t_{0}),p_{1}^{(12)}(t_{0}),p_{2}^{(12)}(t_{0})\right)=(a_{0},1,0,0),$
where $a_{0}\in\mathbb{C}$. This solution is expressed in terms of formal power series as
$\begin{split}&q_{1}^{(1)}=a_{0}+(t-t_{0})^{\lambda_{a}}\sum_{k=1}^{\infty}a_{k}(t-t_{0})^{k},\quad q_{2}^{(1)}=1+(t-t_{0})^{\lambda_{b}}\sum_{k=1}^{\infty}b_{k}(t-t_{0})^{k},\\ &p_{1}^{(12)}=(t-t_{0})^{\lambda_{c}}\sum_{k=1}^{\infty}c_{k}(t-t_{0})^{k},\quad p_{2}^{(12)}=(t-t_{0})^{\lambda_{d}}\sum_{k=1}^{\infty}d_{k}(t-t_{0})^{k},\end{split}$ (5.14)
@@
Substituting solution (5.14) to the fourth equation of system (5.13), we obtain
$t_{0}(t_{0}-1)(\lambda_{d}+1)d_{1}(t-t_{0})^{\lambda_{d}}+O\left((t-t_{0})^{\lambda_{d}+1}\right)=-t_{0}+1+o(1).$
Here we denote the function $f(t)$ satisfying $\lim_{t\to t_{0}}\left|\frac{f(t)}{(t-t_{0})^{k}}\right|=0$ by $o\left((t-t_{0})^{k}\right)$. It follows that
$\lambda_{d}=0,\quad d_{1}=-\frac{1}{t_{0}}.$ (5.15)
@@
$\lambda_{b}=0,\quad b_{1}=-\frac{\alpha_{2}}{t_{0}}$ (5.16)
from the second equation of system (5.13),
$\lambda_{a}=0$ (5.17)
from the first equation and
$\lambda_{c}=0$ (5.18)
from the third equation. Under conditions (5.15), (5.16), (5.17) and (5.18), we also obtain
$a_{1}=-\frac{2a_{0}(a_{0}-1)(t_{0}a_{0}-1)}{t_{0}-1}c_{1}+A_{1},$
where $A_{1}\in\mathbb{C}(t_{0})[\alpha_{0},\ldots,\alpha_{5},\eta,a_{0}]$ from the second equation. Then we see that $c_{1}$ turns to be an arbitrary complex number and $a_{1}$ is determined uniquely by $c_{1}$. Moreover, in a similar manner, $c_{2}$ and $d_{2}$ are determined uniquely as elements of $\mathbb{C}(t_{0})[\alpha_{0},\ldots,\alpha_{5},\eta,a_{0},c_{1}]$ and we obtain
$a_{2}=-\frac{(a_{0}-1)(t_{0}a_{0}+1)}{2(t_{0}-1)}b_{2}+A_{2},$
where $A_{2}\in\mathbb{C}(t_{0})[\alpha_{0},\ldots,\alpha_{5},\eta,a_{0},c_{1}]$. Then we see that $b_{2}$ turns to be an arbitrary complex number and $a_{2}$ is determined uniquely by $b_{2}$.
@@
$\frac{q_{2}^{(1)}-1}{p_{2}^{(12)}}$ $=\frac{-\alpha_{2}t_{0}^{-1}(t-t_{0})+b_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right)}{-t_{0}^{-1}(t-t_{0})+d_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right)}$
$=\frac{-\alpha_{2}t_{0}^{-1}+b_{2}(t-t_{0})+O\left((t-t_{0})^{2}\right)}{-t_{0}^{-1}+d_{2}(t-t_{0})+O\left((t-t_{0})^{2}\right)}$
$=\alpha_{2}+\frac{-t_{0}^{-1}b_{2}+\alpha_{2}t_{0}^{-1}d_{2}}{t_{0}^{-2}}(t-t_{0})+O\left((t-t_{0})^{2}\right)$
$=\alpha_{2}+t_{0}(-b_{2}+\alpha_{2}d_{2})(t-t_{0})+O\left((t-t_{0})^{2}\right),$
@@
$\frac{\alpha_{2}-\frac{q_{2}^{(1)}-1}{p_{2}^{(12)}}}{p_{2}^{(12)}}$ $=\frac{t_{0}(b_{2}-\alpha_{2}d_{2})(t-t_{0})+O\left((t-t_{0})^{2}\right)}{-t_{0}^{-1}(t-t_{0})+O\left((t-t_{0})^{2}\right)}$
$=\widetilde{b}_{0}+O(t-t_{0}),$
where $\widetilde{b}_{0}=-t_{0}^{2}(b_{2}-\alpha_{2}d_{2})$. We also have
$\frac{p_{1}^{(12)}}{p_{2}^{(12)}}=\widetilde{c}_{0}+O(t-t_{0}),$
where $\widetilde{c}_{0}=-t_{0}c_{1}$ in a similar manner. Set
$x_{2}^{[12]}=q_{1}^{(1)},\quad y_{2}^{[12]}=\frac{\alpha_{2}-\frac{q_{2}^{(1)}-1}{p_{2}^{(12)}}}{p_{2}^{(12)}},\quad z_{2}^{[12]}=\frac{p_{1}^{(12)}}{p_{2}^{(12)}},\quad w_{2}^{[12]}=p_{2}^{(12)}.$
@@
$\left(x_{2}^{[12]}(t_{0}),y_{2}^{[12]}(t_{0}),z_{2}^{[12]}(t_{0}),w_{2}^{[12]}(t_{0})\right)=\left(a_{0},\widetilde{b}_{0},\widetilde{c}_{0},0\right).$
In the last, we consider the case that the coefficients $\widetilde{b}_{0}$ and $\widetilde{c}_{0}$ tend to infinity. Set
$x=q_{1}^{(1)},\quad y=-\frac{\frac{q_{2}^{(1)}-1}{p_{1}^{(12)}}-\alpha_{2}\frac{p_{2}^{(12)}}{p_{1}^{(12)}}}{p_{2}^{(12)}},\quad z=p_{1}^{(12)},\quad w=\frac{p_{2}^{(12)}}{p_{1}^{(12)}}$
and
$x^{\prime}=q_{1}^{(1)},\quad y^{\prime}=\alpha_{2}-\frac{q_{2}^{(1)}-1}{p_{2}^{(12)}},\quad z^{\prime}=\frac{p_{1}^{(12)}}{\alpha_{2}-\frac{q_{2}^{(1)}-1}{p_{2}^{(12)}}},\quad w^{\prime}=\frac{p_{2}^{(12)}}{\alpha_{2}-\frac{q_{2}^{(1)}-1}{p_{2}^{(12)}}}.$
By using solution (5.14) with (5.15), (5.16), (5.17) and (5.18), we have
$x=a_{0}+O(t-t_{0}),\quad y=\frac{\widetilde{b}_{0}}{\widetilde{c}_{0}}+O(t-t_{0}),$
$z=O(t-t_{0}),\quad w=\frac{1}{\widetilde{c}_{0}}+O(t-t_{0})$
and
$x^{\prime}=a_{0}+O(t-t_{0}),\quad y^{\prime}=O(t-t_{0}),$
$z^{\prime}=\frac{\widetilde{c}_{0}}{\widetilde{b}_{0}}+O(t-t_{0}),\quad w^{\prime}=\frac{1}{\widetilde{b}_{0}}+O(t-t_{0}).$
Hence we obtain
$(x(t_{0}),y(t_{0}),z(t_{0}),w(t_{0}))=\left(a_{0},\frac{\widetilde{b}_{0}}{\widetilde{c}_{0}},0,\frac{1}{\widetilde{c}_{0}}\right)$
and
$(x^{\prime}(t_{0}),y^{\prime}(t_{0}),z^{\prime}(t_{0}),w^{\prime}(t_{0}))=\left(a_{0},0,\frac{\widetilde{c}_{0}}{\widetilde{b}_{0}},\frac{1}{\widetilde{b}_{0}}\right).$
On the other hand, the systems on new coordinates are described as
$t(t-1)wdx-Xdt=0,$
$t(t-1)wdy-Ydt=0,$
$t(t-1)wdz-Zdt=0,$
$t(t-1)dw-Wdt=0,$
where $X,Y,Z,W\in\mathbb{C}[x,y,z,w,t]$ and
$t(t-1)w^{\prime}dx^{\prime}-X^{\prime}dt=0,$
$t(t-1)w^{\prime}dy^{\prime}-Y^{\prime}dt=0,$
$t(t-1)w^{\prime}dz^{\prime}-Z^{\prime}dt=0,$
$t(t-1)dw^{\prime}-W^{\prime}dt=0,$
where $X^{\prime},Y^{\prime},Z^{\prime},W^{\prime}\in\mathbb{C}\left[x^{\prime},y^{\prime},z^{\prime},w^{\prime},t\right]$. Then all of the leaves corresponding to the solutions to these systems with the initial conditions
$\left(x(t_{0}),y(t_{0}),z(t_{0}),w(t_{0})\right)=\left(a_{0},\widehat{b}_{0},0,0\right),$
where $\widehat{b}_{0}\in\mathbb{C}$ and
$\left(x^{\prime}(t_{0}),y^{\prime}(t_{0}),z^{\prime}(t_{0}),w^{\prime}(t_{0})\right)=(a_{0},0,\widehat{c}_{0},0),$
where $\widehat{c}_{0}\in\mathbb{C}$ are both vertical leaves. We do not state their detail here.
6 Resolving the accessible singularities $P^{0}$
It is not suitable to use the blowing up method to resolve the accessible singularity $P^{0}$, because the explicit formula of the system after blowing up becomes more complicated. Therefore we use the method in terms of formal power series and the symmetry of system (2.2) in this section. We give the result of this section in advance.